OpenStudy (anonymous):

3√ 2-2√ 3, can this be simplified

OpenStudy (swissgirl):

Nope it cant. If we had the same number inside both square roots then we can but in this case they are diff one is 2 and the other is 3

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

by any chance do you know how to simp this? 2√ 3+√ 27-√ 1/3

OpenStudy (swissgirl):

hmmmm is it like this? \(\large 2\sqrt{3} + \sqrt{27} - \sqrt{1 \over 3} \) \(\large2\sqrt{3}+\sqrt{3*3*3}-\sqrt{1 \over 3} \) \(\large2\sqrt{3}+3\sqrt{3}-\sqrt{1 \over 3} \) \(\large 5\sqrt{3}-\sqrt{1 \over 3}\)